# for Hand Motion Generation in Piano Performance

### v1

@@ Abstract
Recently, artificial intelligence techniques for education have been received increasing attentions, while it still remains an open problem to design the effective music instrument instructing systems. Although key presses can be directly derived from sheet music, the transitional movements among key presses require more extensive guidance in piano performance. In this work, we construct a piano-hand motion generation benchmark to guide hand movements and fingerings for piano playing. To this end, we collect an annotated dataset, PianoMotion10M, consisting of 116 hours of piano playing videos from a bird’s-eye view with 10 million annotated hand poses. We also introduce a powerful baseline model that generates hand motions from piano audios through a position predictor and a position-guided gesture generator. Furthermore, a series of evaluation metrics are designed to assess the performance of the baseline model, including motion similarity, smoothness, positional accuracy of left and right hands, and overall fidelity of movement distribution. Despite that piano key presses with respect to music scores or audios are already accessible, PianoMotion10M aims to provide guidance on piano fingering for instruction purposes. The dataset and source code can be accessed at https://agnjason.github.io/PianoMotion-page.
@@ 1 Introduction
The process of learning has been significantly improved with artificial intelligence techniques, which enable individuals to enhance their skills under the guidance of an AI coach [64, 66]. This can be extended into learning to play the musical instruments. Particularly, piano performance requires a profound understanding of the underlying relationship between musical compositions and their corresponding physical motions. As the rigorous practice and training program are necessary for athletes to master a diverse range of expressive human poses, it entails extensive practices to achieve proficiency in piano fingering and hand movement. To facilitate access to playing guidance, the development of AI piano coach has been spurred.
Large-scale piano-motion datasets are the foundation of a nuanced approach for motion generation, which offer valuable guidance for physical performance and musical expression. The computational challenge of motion generation lies in capturing the nonlinear relationship between musical pieces and the intricate hand motions required for piano playing. The hand poses vary for the same note across different melodies. Moreover, the dynamic nature of musical expression demands a level of continuous motion, which is challenging to be learnt from small datasets. These limitations underscore the urgent need for a large-scale piano-motion dataset.
To address the deficiency of the dataset for guiding hand movements and fingerings in playing piano, we introduce a large-scale 3D piano-motion dataset named PianoMotion10M. As shown in Fig. 1, PianoMotion10M contains piano audio tracks, Musical Instrument Digital Interface (MIDI) files, and annotated hand motions with their corresponding videos, meticulously collected from the Internet. It comprises 1,966 pairs of video and music, with a total duration of 116 hours and 10 million annotated frames. The parametric MANO hand model [47] is employed to represent hand gestures. Our constructed dataset offers a diverse range of music styles and piano techniques, which addresses the demands of various preferences and skill levels.
@@
Traditional applications like PianoPlayer [45] are adept at generating static hand gestures and positions for piano scores, which typically make use of classifiers to predict the proper fingering combinations. However, they often fail to capture the diversity and continuity inherent of the piano performance in PianoMotion10M. Both rule-based methods [5, 33] and HMM-based approaches [69, 41] can estimate fingerings while they primarily focus on the local fingering constraints of continuous notes. Consequently, they often overlook crucial information like long-range fingering relationships, while our task aims to estimate the motions of long clips.
To address these limitations, a novel baseline model is presented to show the effectiveness of PianoMotion10M, which is able to generate realistic hand motions from piano melodies. Given a piece of piano music, our model can locate the positions of both hands and generate a long sequence of hand gestures for the performance. It effectively learns the music-position correlation through an efficient position predictor and produces continuous gestures with a position-guided gesture generator based on a diffusion probabilistic model. To assess our baseline model, we propose several evaluation metrics, including Frechet Gesture Distance and Wasserstein Gesture Distance to measure the fidelity of each hand motions, Frechet Inception Distance with a pre-trained auto-encoder to investigate the motion quality of double hands, and Position Distance to assess the accuracy of hand positioning, and Smoothness of the generated motions.
@@ 2 Related Work
Motion-music Datasets. While multi-modal datasets [52, 55, 36, 31] become the key of various learning tasks, there remains a significant gap in the availability of repository specially designed for music-conditional motion generation. Although the existing hand gesture datasets [39, 38, 30, 16, 67] contain a large number of image-hand gesture pairs, they do not have audio or other related information. This limitation hinders them from the generative tasks, since they mainly focus on hand reconstruction and pose estimation. Recently, datasets like AIST++ [32] and TikTok [77] are tailored for music-dance learning, which provide limited music segments, typically less than 5 hours in duration. Moryossef et al. [40] manage to automatically detect which fingers pressed the key of the piano and provide a dataset of piano-fingering. However, it does not provide continuous hand gesture movements. Therefore, there is a crucial need to build a piano-motion dataset specially tailored for motion generation tasks. To this end, we introduce the PianoMotion10M dataset, which comprises extensive piano music and their corresponding hand motion annotations.
@@
3D Human Motion Synthesis. The problem of generating realistic and controllable 3D human motion sequences has been a subject of long-standing study. By taking advantage of 2D keypoint detection [8], the synthesis of 2D skeletons has been extensively explored [46, 53, 15]. Considerable research efforts have been devoted to 2D speech-driven head generation for facial mouth and lip motion generation [72, 10, 19, 25, 21], which usually employ either image-driven or voice-driven methods to produce realistic videos of speaking individuals. However, the expressive capabilities of 2D pose skeletons are limited and they are not applicable to 3D character models. Recent methods for full-body 3D dance generation have utilized LSTM [58, 68, 78], GANs [57, 17] or transformer encoders [32, 24, 54]. To generate vivid talking head videos, extensive research has been conducted in the field of speech-driven 3D facial animation [7, 13, 60, 14, 73]. EmoTalk [44] animates emotional 3D faces from speech input by generating controllable personal and emotional styles. Tian et al. [61] introduce a speed controller and a face region controller to enhance stability during the head generation process.
@@
In the domain of hand motion generation, it has been primarily categorized into rule-based methods [9, 23, 56] and data-driven approaches [29, 12, 34, 70, 6]. For instance, Speech2Gesture [17] utilizes conditional generative adversarial networks to generate personalized 2D keypoints from audio. Ahuja et al. [1] propose a method for personalized motion transfer. Ao et al. [62] learn the mapping between the speech and gestures from data using a combined network structure of the vector quantized variational auto-encoder (VQ-VAE). Beyond 2D keypoints generation, TriModal [70] extracts different upper body movements from TED talks and designs a LSTM-based neural network conditioned on audio, text, and identity to generate co-speech gestures.
@@
Recently, diffusion models have achieved promising results in generating human motions. Previous works such as MDM [59] and MotionDiffuse [74] have produced realistic motion inspired by denoising diffusion probabilistic models (DDPM) [20]. PhysDiff [71] extends MDM by imposing physical constraints. MLD [11] utilizes latent carrier DDPM for forward denoising and reverse diffusion in motion latent space. MAA [3] enhances the performance of non-distributed data by pre-training diffusion models. Zhang et al. [75] introduce retrieval-enhanced DDPM, which improves the text-to-motion functionality in distribution.
@@ 3 PianoMotion10M Dataset
It is a formidably challenging task to map piano music to hand motions due to the significant influence of note sequences and positions on hand movements and fingering. Lacking diverse data may lead to the inferior performance on estimating various hand poses in piano playing. To capture the variability, we present the first large-scale piano-motion dataset, PianoMotion10M, which comprises 1,966 piano performance videos along with 10 million hand poses and their corresponding MIDI files, resulting in an overall duration of 116 hours. Fig. 2 presents an example of our dataset. These videos are segmented into 16,739 individual clips with a length of 30 seconds. To ensure comprehensive evaluation, we extract 7,519 clips for training, 821 for validation and 8,399 for testing. The detailed comparisons with the existing datasets on hands and 3D human motions are summarized into Tab. 1.
Table 1: Comparison between different hand and motion datasets. The proposed PianoMotion10M dataset consists of piano music with corresponding hand poses for hand motion generation. Existing hand-image datasets are listed in the first four rows, and music-motion datasets are presented in the subsequent four rows for reference.
  Dataset Pose Size Subject Music MIDI Duration(hour)
FreiHAND [79] 134K 32 -
InterHand2.6M [39] 2.6M 27 -
RGB2Hands [65] 1K 2 -
Re:InterHand [37] 1.5M 10 -
GrooveNet [2] - 1 0.38
DanceNet [78] - 2 0.96
EA-MUD [57] - - 0.35
AIST++ [32] 10.1M 30 5.19
PianoMotion10M 10.5M 14 116.16
@@ 3.1 Data Collection
There is a wealth of videos and live streams dedicated to musical instrument performances and tutorials from the Internet. Note that each individual has a unique playing style, we firstly select $14$ piano experts from Bilibili [^1] https://www.bilibili.com, one of the most popular video-sharing platforms in China. and collect a total of 3,647 candidate videos. To ensure consistency and enhance the quality of our dataset, we conduct pre-processing with five pivotal factors, including resolution, audio quality, camera perspective, presence of multiple individuals, and visibility of hands. We manually select pure piano music to ensure that the audio contained no human vocals or sounds from other instruments. Moreover, videos with a resolution lower than $1080\times 1920$ are discarded to ensure the quality of our dataset. To enhance the observation of hand movements and gestures, we choose videos with a bird’s-eye view to minimize hand occlusion. Furthermore, we remove those videos where hand regions are frequently obstructed or invisible during performance. Additionally, we do not take consideration of videos with multiple piano players, which is beyond the scope of this dataset. By addressing these factors, the overall quality and coherence of the dataset have been substantially enhanced. Following this preprocessing stage, the result is a collection of $1,966$ high-quality raw videos. Each one showcases the piano playing by an individual, which is captured from a bird’s-eye view along with pure piano audio.
@@ 3.2 Data Annotations
MIDI is a universal digital protocol to store musical data for various musical devices, which is served as a digital representation of musical notes, volume, tempo, and other performance parameters. Once video and pure piano audio pairs are collected, we transcribe the piano performances into MIDI files by utilizing a state-of-the-art automatic piano transcription [28]. To ensure the accuracy of the conversion results, we replay MIDI files and compare them with the original music tracks. Those files with high discrepancies are adjusted.
@@
Hand Pose is captured via the parametric hand model MANO [47], which serves as our hand prior model. It effectively maps the pose parameter $\theta\in\mathbb{R}^{J\times 3}$ with $J$ per-bone parts and the shape parameter $\rho\in\mathbb{R}^{10}$ onto a template mesh $\bar{\mathcal{M}}$ with vertices $V$. The MANO model enables the simulation of various hand gestures and movements during piano performances, which provides an effective way to study the relationship between hand gestures in playing piano. Due to the non-uniformity of hand sizes and positions in the collected videos, we first employ the MediaPipe hand detection framework [35] to obtain bounding box of the hand region. Video frames are cropped according to the detected hand bounding boxes to enhance the robustness of the results. To this end, hand poses in collected videos are annotated using HaMeR [42]. HaMeR follows a fully transformer-based architecture and reconstructs hand models with increased accuracy and robustness. All images having hands detected by MediaPipe are annotated with hand poses by HaMeR, which result in a dataset of 10 million image-pose pairs.
@@
To enhance the smoothness and continuity of our dataset, the generated hand poses require to be cleaned and refined. The results of MediaPipe and HaMeR are usually accurate in most cases, while some inferior results may occur due to rapid motion and image blurring. The Hampel filter [43] with a window size of $20$ is utilized to identify these outliers. The outliers and the timestamps where hands are undetected are firstly labeled as missing values. Within a small period, hand movements can be considered as the motion with constant speed. Therefore, the small gap can be interpolated bilaterally. To address these missing data in the time series, missing segments with the frame length $\delta$ less than $30$ frames are filled by linear interpolation with respect to their surrounding values. The others are considered as periods when the hand is invisible. To ensure the reliability of detected hands, a similar strategy is employed to label observations with excessively short duration ($\delta<15$) as invisible. Finally, to ensure the smoothness of hand motions, we make use of a Savitzky-Golay filter [51] for data smoothing. The annotated hand poses are manually checked to ensure their quality.

### v2

@@ Abstract
Recently, artificial intelligence techniques for education have been received increasing attentions, while it still remains an open problem to design the effective music instrument instructing systems. Although key presses can be directly derived from sheet music, the transitional movements among key presses require more extensive guidance in piano performance. In this work, we construct a piano-hand motion generation benchmark to guide hand movements and fingerings for piano playing. To this end, we collect an annotated dataset, PianoMotion10M, consisting of 116 hours of piano playing videos from a bird’s-eye view with 10 million annotated hand poses. We also introduce a powerful baseline model that generates hand motions from piano audios through a position predictor and a position-guided gesture generator. Furthermore, a series of evaluation metrics are designed to assess the performance of the baseline model, including motion similarity, smoothness, positional accuracy of left and right hands, and overall fidelity of movement distribution. Despite that piano key presses with respect to music scores or audios are already accessible, PianoMotion10M aims to provide guidance on piano fingering for instruction purposes. The source code and dataset can be accessed at https://github.com/agnJason/PianoMotion10M.
@@ 1 Introduction
The process of learning has been significantly improved with artificial intelligence techniques, which enable individuals to enhance their skills under the guidance of an AI coach (Wang et al., 2019; Zakka et al., 2023; Wang et al., 2024). This can be extended into learning to play the musical instruments. Particularly, piano performance requires a profound understanding of the underlying relationship between musical compositions and their corresponding physical motions. As the rigorous practice and training program are necessary for athletes to master a diverse range of expressive human poses, it entails extensive practices to achieve proficiency in piano fingering and hand movement. To facilitate access to playing guidance, the development of AI piano coach has been spurred.
Large-scale piano-motion datasets are the foundation of a nuanced approach for motion generation, which offer valuable guidance for physical performance and musical expression. The computational challenge of motion generation lies in capturing the nonlinear relationship between musical pieces and the intricate hand motions required for piano playing. The hand poses vary for the same note across different melodies. Moreover, the dynamic nature of musical expression demands a level of continuous motion, which is challenging to be learned from small datasets. These limitations underscore the urgent need for a large-scale piano-motion dataset.
To address the deficiency of the dataset for guiding hand movements and fingerings in playing piano, we introduce a large-scale 3D piano-motion dataset named PianoMotion10M. As shown in Fig. 1, PianoMotion10M contains piano audio tracks, Musical Instrument Digital Interface (MIDI) files, and annotated hand motions with their corresponding videos, meticulously collected from the Internet. It comprises 1,966 pairs of video and music, with a total duration of 116 hours and 10 million annotated frames. The parametric MANO hand model (Romero et al., 2017) is employed to represent hand gestures. Our constructed dataset offers a diverse range of music styles and piano techniques, which addresses the demands of various preferences and skill levels.
@@
Traditional applications like PianoPlayer (PianoPlayer, 2018) are adept at generating static hand gestures and positions for piano scores, which typically make use of classifiers to predict the proper fingering combinations. However, they often fail to capture the diversity and continuity inherent of the piano performance in PianoMotion10M. Both rule-based methods (Balliauw et al., 2017; Lin & Liu, 2006) and HMM-based approaches (Yonebayashi et al., 2007; Nakamura & Sagayama, 2015) can estimate fingerings while they primarily focus on the local fingering constraints of continuous notes. Consequently, they often overlook crucial information like long-range fingering relationships, while our task aims to estimate the motions of long clips.
To address these limitations, a novel baseline model is presented to show the effectiveness of PianoMotion10M, which is able to generate realistic hand motions from piano melodies. Given a piece of piano music, our model can locate the positions of both hands and generate a long sequence of hand gestures for the performance. It effectively learns the music-position correlation through an efficient position predictor and produces continuous gestures with a position-guided gesture generator based on a diffusion probabilistic model. To assess our baseline model, we propose several evaluation metrics, including Fréchet Gesture Distance and Wasserstein Gesture Distance to measure the fidelity of each hand motions, Fréchet Inception Distance with a pre-trained auto-encoder to investigate the motion quality of double hands, and Position Distance to assess the accuracy of hand positioning, and Smoothness of the generated motions.
@@ 2 Related Work
Motion-music Datasets. While multi-modal datasets (Schuhmann et al., 2021; Srinivasan et al., 2021; Miech et al., 2019; Lee et al., 2021) become the key of various learning tasks, there remains a significant gap in the availability of repository specially designed for music-conditional motion generation. Although the existing hand gesture datasets (Moon et al., 2020b; a; Kwon et al., 2021; Gan et al., 2024; Wu et al., 2023) contain a large number of image-hand gesture pairs, they do not have audio or other related information. This limitation hinders them from the generative tasks, since they mainly focus on hand reconstruction and pose estimation. Recently, datasets like AIST++ (Li et al., 2021) and TikTok (Zhu et al., 2022) are tailored for music-dance learning, which provide limited music segments, typically less than 5 hours in duration. Moryossef et al. (Moryossef et al., 2023) manage to automatically detect which fingers pressed the key of the piano and provide a dataset of piano-fingering. However, it does not provide continuous hand gesture movements. Therefore, there is a crucial need to build a piano-motion dataset specially tailored for motion generation tasks. To this end, we introduce the PianoMotion10M dataset, which comprises extensive piano music and their corresponding hand motion annotations.
Piano Hand Datasets. Yamamoto et al. (Yamamoto et al., 2010) employed a key-frame technique to generate natural hand motions for piano playing from inputted music scores. Liang et al. (Liang et al., 2016) used an RGB-D camera to collect 7,200 depth images and designed models to estimate fingertip movements and predict finger tappings. The CMU Panoptic Studio dataset (Joo et al., 2015; 2017; Simon et al., 2017) provided a multi-view dataset of human-object interactions, including annotated hand and face postures. Simon et al. (Simon et al., 2017) introduced a real-time hand keypoint detector and a markerless 3D hand motion capture system capable of reconstructing challenging hand-object interactions and musical performances. TwohandsMusic (Seo et al., 2019) presented two shallow networks to estimate 3D hand poses and tap gestures, along with a dataset of 85,000 images of hand movements while playing the piano. To help beginners improve their playing techniques, Johnson et al. (Johnson et al., 2020) discussed automatic assessment of pianists’ hand postures, which categorizes postures as correct posture, low wrists, or flat hands. Kun et al. (Su et al., 2020) collected piano performance videos to generate the music for videos directly. Moryossef et al. (Moryossef et al., 2023) managed to automatically detect which fingers pressed the key of the piano and provided a dataset of piano-fingering. Wu et al. (Wu et al., 2023) proposed a marker-removal approach for collecting bare-hand data, which included a precise ground truth alongside a large-scale pianist 3D hand dataset, PianoHand2.5M. This dataset provided advanced annotated images for estimating hand postures in marked-hand images.
Our dataset and baseline primarily focus on researching the relationships between piano music and hand postures. More comparisons with other existing hand datasets are shown in the attached pdf document. Since most of existing piano-hand datasets are not publicly available, our open-source release of PianoMotion10M will be of substantial importance to advancing research in this field.
@@
3D Human Motion Synthesis. The problem of generating realistic and controllable 3D human motion sequences has been a subject of long-standing study. By taking advantage of 2D keypoint detection (Cao et al., 2017), the synthesis of 2D skeletons has been extensively explored (Ren et al., 2020; Shiratori et al., 2006; Ferreira et al., 2021). Considerable research efforts have been devoted to 2D speech-driven head generation for facial mouth and lip motion generation (Zakharov et al., 2019; Chen et al., 2019; Guo et al., 2021; Ji et al., 2021; Hong et al., 2022), which usually employ either image-driven or voice-driven methods to produce realistic videos of speaking individuals. However, the expressive capabilities of 2D pose skeletons are limited and they are not applicable to 3D character models. Recent methods for full-body 3D dance generation have utilized LSTM (Tang et al., 2018; Yalta et al., 2019; Zhuang et al., 2022), GANs (Sun et al., 2020; Ginosar et al., 2019) or transformer encoders (Li et al., 2021; Huang et al., 2020; Siyao et al., 2022). To generate vivid talking head videos, extensive research has been conducted in the field of speech-driven 3D facial animation (Cao et al., 2005; Cudeiro et al., 2019; Tian et al., 2019; Fan et al., 2022; Zhang et al., 2021). EmoTalk (Peng et al., 2023) animates emotional 3D faces from speech input by generating controllable personal and emotional styles. Tian et al. (Tian et al., 2024) introduce a speed controller and a face region controller to enhance stability during the head generation process.
@@
In the domain of hand motion generation, it has been primarily categorized into rule-based methods (Cassell et al., 2001; Huang & Mutlu, 2012; Starke et al., 2022) and data-driven approaches (Kopp et al., 2006; Chiu et al., 2015; Liu et al., 2022; Yoon et al., 2020; Bhattacharya et al., 2021). For instance, Speech2Gesture (Ginosar et al., 2019) utilizes conditional generative adversarial networks to generate personalized 2D keypoints from audio. Ahuja et al. (Ahuja et al., 2020) propose a method for personalized motion transfer. Ao et al. (Van Den Oord et al., 2017) learn the mapping between the speech and gestures from data using a combined network structure of the vector quantized variational auto-encoder (VQ-VAE). Beyond 2D keypoints generation, TriModal (Yoon et al., 2020) extracts different upper body movements from TED talks and designs a LSTM-based neural network conditioned on audio, text, and identity to generate co-speech gestures.
@@
Recently, diffusion models have achieved promising results in generating human motions. Previous works such as MDM (Tevet et al., 2022) and MotionDiffuse (Zhang et al., 2024) have produced realistic motion inspired by denoising diffusion probabilistic models (DDPM) (Ho et al., 2020). PhysDiff (Yuan et al., 2023) extends MDM by imposing physical constraints. MLD (Chen et al., 2023) utilizes latent carrier DDPM for forward denoising and reverse diffusion in motion latent space. MAA (Azadi et al., 2023) enhances the performance of non-distributed data by pre-training diffusion models. Zhang et al. (Zhang et al., 2023) introduce retrieval-enhanced DDPM, which improves the text-to-motion functionality in distribution.
@@ 3 PianoMotion10M Dataset
It is a formidably challenging task to map piano music to hand motions due to the significant influence of note sequences and positions on hand movements and fingering. Lacking diverse data may lead to the inferior performance on estimating various hand poses in piano playing. To capture the variability, we present the first large-scale piano-motion dataset, PianoMotion10M, which comprises 1,966 piano performance videos along with 10 million hand poses and their corresponding MIDI files, resulting in an overall duration of 116 hours. Fig. 2 presents an example of our dataset. These videos are segmented into 16,739 individual clips with a length of 30 seconds. To ensure comprehensive evaluation, we extract 7,519 clips for training, 821 for validation and 8,399 for testing. The dataset is divided based on the original videos to avoid the overlap in piano performances. The detailed comparisons with the existing datasets on hands and 3D human motions are summarized into Tab. 1.
Table 1: Comparison between different hand and motion datasets. The proposed PianoMotion10M dataset consists of piano music with corresponding hand poses for hand motion generation. For reference, the table is organized as follows: the first five rows list existing hand-image datasets, the subsequent four rows present music-motion datasets, and the following four rows display various piano-hand pose datasets. The final row features our dataset.
  Dataset Year Pose Size Subject Music MIDI Duration(h) Available
CMU Panoptic 2017 31K 8 -
FreiHAND 2019 134K 32 -
InterHand2.6M 2020b 2.6M 27 -
RGB2Hands 2020 1K 2 -
Re:InterHand 2023 1.5M 10 -
GrooveNet 2017 - 1 0.38
DanceNet 2022 - 2 0.96
EA-MUD 2020 - - 0.35
AIST++ 2021 10.1M 30 5.19
Barehanded Music 2016 7.2K 2 -
TwohandsMusic 2019 85K 5 -
Piano-fingering 2023 155K - -
PianoHand2.5M 2023 2.5M 21 4.3
PianoMotion10M 2024 10.5M 14 116
@@ 3.1 Data Collection
There is a wealth of videos and live streams dedicated to musical instrument performances and tutorials from the Internet. Note that each individual has a unique playing style, we firstly select $14$ piano experts from Bilibili [^1] https://www.bilibili.com, one of the most popular video-sharing platforms in China. and collect a total of 3,647 candidate videos. To ensure consistency and enhance the quality of our dataset, we conduct pre-processing with five pivotal factors, including resolution, audio quality, camera perspective, presence of multiple individuals, and visibility of hands. Accurate, noise-free audio is essential to our dataset. We manually select pure piano music to ensure the audio contains no human vocals or sounds from other instruments, rather than employing on automated tools Kong et al. (2020). Moreover, videos with a resolution lower than $1080\times 1920$ are discarded to ensure the quality of our dataset. To enhance the observation of hand movements and gestures, we choose videos with a bird’s-eye view to minimize hand occlusion. Furthermore, we remove those videos where hand regions are frequently (more than 20% frames) obstructed or invisible during performance. Additionally, we do not take consideration of videos with multiple piano players, which is beyond the scope of this dataset. By addressing these factors, the overall quality and coherence of the dataset have been substantially enhanced. Following this preprocessing stage, the result is a collection of $1,966$ high-quality raw videos. Each one showcases the piano playing by an individual, which is captured from a bird’s-eye view along with pure piano audio.
@@ 3.2 Data Annotations
MIDI is a universal digital protocol to store musical data for various musical devices, which is served as a digital representation of musical notes, volume, tempo, and other performance parameters. Some creators provide ground-truth MIDI files for their performances. For the remaining piano performances, automatic piano transcription methods (Kong et al., 2021; Maman & Bermano, 2022; Gardner et al., 2021; Toyama et al., 2023; Cheuk et al., 2023) are employed to convert them into MIDI files, while  (Kong et al., 2021) demonstrates more stable and robust performance in wild data. To ensure the accuracy of the conversion results, we replay MIDI files and compare them with the original music tracks. Files are adjusted if they exceed thresholds such as timing differences over 30 ms, dynamic variations beyond 10%, or noticeable pitch mismatches that alter melody or harmony.
@@
Hand Pose is captured via the parametric hand model MANO (Romero et al., 2017), which serves as our hand prior model. It effectively maps the pose parameter $\theta\in\mathbb{R}^{J\times 3}$ with $J$ per-bone parts and the shape parameter $\rho\in\mathbb{R}^{10}$ onto a template mesh $\bar{\mathcal{M}}$ with vertices $V$. The MANO model enables the simulation of various hand gestures and movements during piano performances, which provides an effective way to study the relationship between hand gestures in playing piano. Due to the non-uniformity of hand sizes and positions in the collected videos, we first employ the MediaPipe hand detection framework (Lugaresi et al., 2019) to obtain bounding box of the hand region. Video frames are cropped according to the detected hand bounding boxes to enhance the robustness of the results. To this end, hand poses in collected videos are annotated using HaMeR (Pavlakos et al., 2023). HaMeR follows a fully transformer-based architecture and reconstructs hand models with increased accuracy and robustness. All images having hands detected by MediaPipe are annotated with hand poses by HaMeR, which result in a dataset of 10 million image-pose pairs.
@@
To enhance the smoothness and continuity of our dataset, the generated hand poses require to be cleaned and refined. The results of MediaPipe and HaMeR are usually accurate in most cases, while some inferior results may occur due to rapid motion and image blurring. The Hampel filter (Pearson et al., 2016) with a window size of $20$ is utilized to identify these outliers. The outliers and the timestamps where hands are undetected are firstly labeled as missing values. Within a small period, hand movements can be considered as the motion with constant speed. Therefore, the small gap can be interpolated bilaterally. To address these missing data in the time series, missing segments with the frame length $\delta$ less than $30$ frames are filled by linear interpolation with respect to their surrounding values. The others are considered as periods when the hand is invisible. To ensure the reliability of detected hands, a similar strategy is employed to label observations with excessively short duration ($\delta<15$) as invisible. Finally, to ensure the smoothness of hand motions, we make use of a Savitzky-Golay filter (Schafer, 2011) with an order of 3 and a window size of 11 for data smoothing, which delivers better performance in terms of noise separation, artifacts and baseline drifts. The annotated hand poses are manually checked to ensure their quality.